OpenStudy (anonymous):

Let f(x)= √(x-1). Find all the values of x in the interval (1,5) guaranteed by the mean value theorem

OpenStudy (anonymous):

you need a couple numbers first

OpenStudy (anonymous):

we need to compute \[\frac{f(b)-f(a)}{b-a}\] for \(f(x)=\sqrt{x-1},a=1,b=5\)

OpenStudy (anonymous):

i get \[f(1)=\sqrt{1-1}=0,f(5)=\sqrt{5-1}=\sqrt{4}=2\] and so \[\frac{f(5)-f(1)}{5-1}=\frac{2-0}{4}=\frac{1}{2}\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

k, now the next thing you need is the derivative of \(f(x)=\sqrt{x-1}\) you have that?

OpenStudy (anonymous):

no but i can

OpenStudy (anonymous):

you should do it in your head let me know when you get it and what method you used

OpenStudy (anonymous):

1/2(x-1)^-1/2 chain rule lol

OpenStudy (anonymous):

ok good, but it would be better not to use exponential notation, because you are going to have to evaluate that thing

OpenStudy (anonymous):

but thats how i leaned how to derrive chain rules lol show me what you mean

OpenStudy (anonymous):

here is what i recommend the square root is a very common function, so i would just remember that the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

ok makes sence

OpenStudy (anonymous):

it never changes, so there is no reason to always rewrite the square root as \(x^{\frac{1}{2}}\) and the use the power rule and then convert back

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so your derivative is, by the chain rule \[\frac{1}{2\sqrt{x-1}}\]

OpenStudy (anonymous):

it is what you had in any case, but now it is in a usable form your last job is to solve \[\frac{1}{2\sqrt{x-1}}=\frac{1}{2}\] for \(x\) and you are done

OpenStudy (anonymous):

wait 1/2 isnt one of my answers

OpenStudy (anonymous):

\(\frac{1}{2}\) is not the solution you have to solve this equation \[\frac{1}{2\sqrt{x-1}}=\frac{1}{2}\]

OpenStudy (anonymous):

why did you set it = 1/2?

OpenStudy (anonymous):

the mean value theorem says that if blah blah blah there is a number \(c\in (a,b)\) with \[f'(c)=\frac{f(b)-f(a)}{b-a}\] here we have \[f'(c)=\frac{1}{2\sqrt{c}}\] and \[\frac{f(5)-f(1)}{5-1}=\frac{2-0}{4}=\frac{1}{2}\]

OpenStudy (anonymous):

damn typo i meant \[f'(c)=\frac{1}{2\sqrt{c-1}}\]

OpenStudy (anonymous):

that is why i we set it to \(\frac{1}{2}\) because that is the number we got for \[\frac{f(b)-f(a)}{b-a}\] for \(a=1,b=5,f(x)=\sqrt{x-1}\)

OpenStudy (anonymous):

ok i get but how do i solve for x ?

OpenStudy (anonymous):

i would multiply both sides by 2 first

OpenStudy (anonymous):

can you help me not sure how to start

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{x-1}}=\frac{1}{2}\\ \frac{1}{\sqrt{x-2}}=1\]

OpenStudy (anonymous):

that means \(\sqrt{x-1}=1\) oh another typo sorry ignore the 2

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{x-1}}=\frac{1}{2}\\ \frac{1}{\sqrt{x-1}}=1\\ \sqrt{x-1}=1\] etc

OpenStudy (anonymous):

yeah, 2

OpenStudy (anonymous):

so the answer to my question is b) 2 only ????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you satellite :) i think i've seen you in google lol. I remember satellite and purple. THank you <3

OpenStudy (anonymous):

google? really?

OpenStudy (anonymous):

yup lol

OpenStudy (anonymous):

did you google "satellite73"? i would guess you got pictures of cars

OpenStudy (anonymous):

no it was a math problem and you were the instructor lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok thanks no i must read hamlet for english lol bye

OpenStudy (anonymous):

hamlet more fun than math enjoy